OpenStudy (anonymous):

Find the missing value: Principal $14,200; Rate x; Time 7 years; Simple Interest $19,880 I have to find the rate and the answer I got was 0.2 but that's none of the options given. Please help!

OpenStudy (texaschic101):

I = PRT but since your looking for rate, change it to I/PT = R I/PT = R 19880/(14,200* 7) = R 19880 / 99400 = R 0.2 = R correct, except you did not finish. The rate has to be in a percent. So 0.2 turned to a percent is 20%. Your answer is 20%